OpenStudy (anonymous):

Solve for m: (2m-1)/(3)=5+(m)/(6)

OpenStudy (anonymous):

m= 32/3 or something

OpenStudy (anonymous):

Okay. So first, you want to mutliply both sides by 3. This will give you: \[2m-1=15+3m/2\] Now add 1 to both sides. This gives you: 2m = 16+3m/2 Multiply both sides by 2: This gives you: 4m = 32+3m Subtract 3m from both sides and you have: m = 32 Hope this helps! Visit my website at: http://www.tutorsean.net for free math and computer science videos!

OpenStudy (anonymous):

@ seanwalsh1984 The problem equation is: \[\frac{2m-1}{3}=5+\frac{m}{6}\]The solution for m in the preceding equation is 32/5, not 32. If m is replaced by the value, 32, your solution, the LHS does not equal the RHS.\[\left\{\frac{2m-1}{3},5+\frac{m}{6}\right\}\text{/.} m\to 32\text{ }\to \text{ }\left\{21,\frac{31}{3}\right\} \]